OpenStudy (anonymous):

can anyone here that can teach me derivatives? plz. m going to fail the exams :(

OpenStudy (anonymous):

Is there specific kinds of derivative you don't understand?

OpenStudy (anonymous):

chain rule, product rule, quotient rule, trig?

OpenStudy (anonymous):

Simplest form of derivative is Power Rule \[\text{Derivative} \text{ of} \left[\text{Ax}^b\right]=\text{Abx}^{b-1}\]

OpenStudy (anonymous):

definition?

OpenStudy (anonymous):

\[\frac{d}{dx} f(x) = \lim_{h \rightarrow 0} {f(x + h) - f(x) \over h}\]

OpenStudy (anonymous):

it is the ratio of the small change and dependent variable to that in the dependent variable

OpenStudy (anonymous):

like polpak said

OpenStudy (anonymous):

actually i need step by step guide from basics, is that possible? plz help. thanx in advance, :)

OpenStudy (anonymous):

Do you have example problem?

OpenStudy (anonymous):

well i need a guider that can teach me from basics.. it that possible here??

OpenStudy (anonymous):

y=x^2 refer to my formula above derivative of y is 2x

OpenStudy (anonymous):

http://www.khanacademy.org/#calculus

OpenStudy (anonymous):

thnks for the link. hope it works for me??

OpenStudy (bahrom7893):

polpak i have a real quick question about linear independence.. it will take like 2-3 mins.. can u help me?

OpenStudy (anonymous):

yes

OpenStudy (bahrom7893):

k here: http://www.twiddla.com/564137

OpenStudy (anonymous):

what problem do you have in derivatives